# Detecting Multimedia Generated by Large AI Models: A Survey

### v1

@@ 2 Generation
$\clubsuit$ Video. In the pursuit of high-quality video generation, recent research has also turned to diffusion models. Early work [125, 126] are based on DDPM [66] for video generation. Current research extends text-to-image (T2I) diffusion models to text-to-video (T2V) generation. Meta proposes Make-A-Video [127] extending a diffusion-based T2I model to T2V through spatiotemporally factorized diffusion model. A more advanced model, Emu Video [128], is their latest video generation milestone. Unlike Make-A-Video, which requires five models, Emu Video is a simple unified architecture for video generation and uses just two diffusion models. Given a text prompt, Google’s Imagen Video [129] generates high-resolution videos using a base video generation model and a sequence of interleaved spatial and temporal video super-resolution models. More recently, Stability AI’s Stable Video Diffusion [130] and Runway’s GEN-2 [131] offer latent video diffusion models for video generation, while Pika [132] introduces a platform aimed at broadening creative possibilities in video creation, as shown in Fig. 2 c).
Datasets. Though building upon the success of T2I generation, T2V generation requires temporally smooth and realistic motion that matches the text, besides high-quality visual content, making it still in the nascent stage compared with image generation. Furthermore, video generation is much more computationally costly than image synthesis. To our best knowledge, only one work [101] contributes their generated talking head videos via LAMs, see Table I.
@@
Datasets. Due to comparatively less research attention on the detection of fake audio, LibriSeVoc [103] in Table I is the only public dataset that includes diffusion model-based vocoders (WaveGrad [134] and DiffWave [105]) so far.
@@
$\clubsuit$ Multimodal. Multimodal learning refers to an embodied learning situation that includes learning multiple modalities such as text, image, video, and audio [137]. From a generation perspective, visual generation tasks, such as text-to-image and text-to-video, are regarded as multimodal generation. The generative models for these tasks are trained to learn visual representations and language understanding for visual generation. From the detection aspect, detectors that learn multiple modalities for forgery detection are categorized into multimodal. In this context, we define "multimodal generation" from a detection perspective, referring to frameworks capable of creating multimodal output. Multimodal generation process normally contains foundation models as encoders (e.g., CLIP [138], ViT [139] ) and decoders (e.g., Stable Diffusion [81] ), and a LLM for taking language-like representations from encoders for semantic understanding and produces modality signal tokens for guiding content output, see Fig. 2 e). Most recent work, such as HuggingGPT [140], AudioGPT [141], and NExT-GPT [142], all are based on the language understanding and complex reasoning capabilities of LLMs, and utilizing existing off-the-shelf multimodal encoders and decorders as tools to execute multimodal input and output.
Datasets. The multimodal datasets in Table I contain multi-types of fake media. For example, $DGM^{4}$ [14] and COCOFake [110] includes synthesized text and image.
@@ 3 Detection
3.5.2)  , tnode, l sep=5mm, [$\clubsuit$ Empirical Study  , nnode,l sep=31.2mm, [Coccomini et al. [143],VERITE [144]   , 5mmnode]] [$\clubsuit$ Localization  , nnode,l sep=5mm, [ Frequency-based   , 4mmnode,l sep=5mm, [UFAFormer [145]   , 5mmnode]] [ Spatial-based   , 4mmnode,l sep=5mm, [HAMMER [14], HAMMER++ [14], Wang et al. [146]   , 5mmnode]]] [$\clubsuit$ Interpretability  , nnode,l sep=31.2mm, [Xu et al. [147]   , 5mmnode]] [$\clubsuit$ Generalization  , nnode,l sep=5mm, [ Contrastive Learning   , 4mmnode,l sep=5mm, [LASTED [148]   , 5mmnode]] [ Prompt Tuning   , 4mmnode,l sep=5mm, [AntifakePrompt [149]   , 5mmnode]] ] [$\clubsuit$ Attribution  , nnode,l sep=31.2mm, [De-Fake [150]   , 5mmnode]]] [Pure
3.5.1)  , tnode,l sep=5mm, [$\clubsuit$ Text-image Inconsistency  , nnode,l sep=31.2mm, [DIDAN [13]   , 5mmnode]] [$\clubsuit$ Text-assisted  , nnode,l sep=31.2mm, [Amoroso et al. [151]   , 5mmnode]] ] ] [Audio
3.4), xnode, l sep=5mm, [Pure
3.4.1)  , tnode, l sep=5mm, [$\clubsuit$ Vocoder-based  , nnode,l sep=31.2mm, [DetectVocoder [103]   , 5anode]]] ] [Video
@@
3.3.1)  , tnode, l sep=5mm, [$\clubsuit$ Generalization  , nnode,l sep=31.2mm, [RevisitVideo [152]   , 5vnode]]] ] [Image
@@
3.2.2)  , tnode,l sep=5mm, [$\clubsuit$ Empirical Study  , nnode,l sep=31.2mm, [Corvi et al. [93, 153], Ricker et al. [88], Cocchi et al. [110], Papa et al. [62], Porcile et al. [154], Mandelli et al. [67], Carriere et al. [155]   , 5inode]] [$\clubsuit$ Robustness  , nnode,l sep=5mm, [ Frequency-based   , 4inode,l sep=5mm, [D4 [156]   , 5inode]] [ Spatial-based   , 4inode,l sep=5mm, [GLFF [157], Xu et al. [158], Local Statistics [159]   , 5inode]]] [$\clubsuit$ Localization  , nnode,l sep=5mm, [ Weakly-supervised   , 4inode,l sep=5mm, [Tantaru et al. [160]   , 5inode]] [ Fully-supervised   , 4inode,l sep=5mm, [IFDL [83], PAL [161], TruFor [87]   , 5inode]]] [$\clubsuit$ Interpretability  , nnode,l sep=31.2mm, [Aghasanli et al. [162]   , 5inode]] [$\clubsuit$ Generalization  , nnode,l sep=31.2mm, [Online Detection [163], Ojha et al. [89], Cozzolino et al. [164], LSDA [165], Fingerprintnet [166], FACTOR [167], Dogoulis et al. [168], DNF [169]   , 5inode]] [$\clubsuit$ Attribution and Model Parsing  , nnode,l sep=5mm, [ Model Parsing   , 4inode,l sep=5mm, [Asnani et al. [91]   , 5inode]] [ Attribution   , 4inode,l sep=5mm, [Guarnera et al. [170]   , 5inode]] ]] [Pure
3.2.1)  , tnode, l sep=5mm, [$\clubsuit$ Frequency-based Methods  , nnode,l sep=31.2mm, [Wavelet-Packets [171], AUSOME [172], Xi et al. [173], Synthbuster [70]   , 5inode]] [$\clubsuit$ Spatial-based Methods  , nnode,l sep=31.2mm, [Zhong et al. [94], Nguyen et al. [174], multiLID [175]   , 5inode]] [$\clubsuit$ Diffuser Fingerprints   , nnode,l sep=31.2mm, [DIF [95], DIRE [84], SeDID [176]   , 5inode]] [$\clubsuit$ Physical/Physiological  , nnode,l sep=31.2mm, [Borji [177], Perspective [178], Lighting [179]   , 5inode]] ] ] [Text
@@
3.1.2)  , tnode, l sep=5mm, [$\clubsuit$ Empirical Study  , nnode,l sep=5mm, [ Paraphrase Detection  , 4tnode,l sep=5mm, [Wahle et al. [27], Becker et al. [180]   , 5tnode]] [ Attribution  , 4tnode,l sep=5mm, [Antoun et al. [181]   , 5tnode]] [ Human Evaluation  , 4tnode,l sep=5mm, [ Guo et al. [50]   , 5tnode]] [ Generalization/Robustness  , 4tnode,l sep=5mm, [ChatLog [182], Zero-Shot [183], Training-based [48], Sarvazyan et al. [184], In The Wild [44]   , 5tnode]]] [$\clubsuit$ Robustness  , nnode,l sep=5mm, [ Stylistic/Consistency  , 4tnode,l sep=5mm, [J-Guard [185], PHD  [186]   , 5tnode]] [ Adversarial Learning  , 4tnode,l sep=5mm, [RADAR [187], OUTFOX [32]   , 5tnode]] [ Adversarial Data Augmentation  , 4tnode,l sep=5mm, [Yang et al. [39], Shi et al. [188], MGTBench [189]   , 5tnode]]] [$\clubsuit$ Interpretability  , nnode,l sep=5mm, [ Polish Ratio  , 4tnode,l sep=5mm, [Yang et al. [39]   , 5tnode]] [ Shapley Additive Explanations  , 4tnode,l sep=5mm, [Mitrovic et al. [190], Check-GPT [38]   , 5tnode]] [ P-values  , 4tnode,l sep=5mm, [Kirchenbauer et al. [191]   , 5tnode]] [ N-gram Overlaps  , 4tnode,l sep=5mm, [DNA-GPT [192]   , 5tnode]]] [$\clubsuit$ Generalization  , nnode,l sep=5mm, [ Contrastive Learning  , 4tnode,l sep=5mm, [Conda [193]   , 5tnode]] [ Structured Search  , 4tnode,l sep=5mm, [Ghostbuster [23]   , 5tnode]]] [$\clubsuit$ Attribution  , nnode,l sep=5mm, [ Origin Tracing  , 4tnode,l sep=5mm, [Sniffer [194]   , 5tnode]] [ Style Representation  , 4tnode,l sep=5mm, [Few-Shot [195]   , 5tnode]] [ Perplexity  , 4tnode,l sep=5mm, [LLMDet [196]   , 5tnode]] [ Stylometric/Statistical  , 4tnode,l sep=5mm, [Linguistic Model [29], GPT-who [197]   , 5tnode]] [ Deep-learning Based  , 4tnode,l sep=5mm, [TuringBench [198], Contra-X [199], XLNet-FT [200], TopRoBERTa [201]   , 5tnode]]]] [Pure
3.1.1)  , tnode, l sep=5mm, [$\clubsuit$ Hard-explainable Methods  , nnode,l sep=5mm, [ Positive-Unlabeled  , 4tnode,l sep=5mm, [ MPU [202]   , 5tnode]] [ Efficient Perturbations  , 4tnode,l sep=5mm, [ Efficient-DetectGPT [203], Fast-DetectGPT [204], DetectLLM [205]   , 5tnode]] [ Log Probabilities Curvature  , 4tnode,l sep=5mm, [DetectGPT [206]   , 5tnode]] [ Perplexity  , 4tnode,l sep=5mm, [HowkGPT [207], GPTZero [208]   , 5tnode]] ] [$\clubsuit$ Easy-explainable Methods  , nnode,l sep=5mm, [ Stylometry/Coherence  , 4tnode,l sep=5mm, [Stylometric Signals [209], Coherence Consistency [210]   , 5tnode]] [ Artifacts  , 4tnode,l sep=5mm, [Pu et al. [211]   , 5tnode]] [ Watermarking  , 4tnode,l sep=5mm, [Distillation-Resistant [212], Fernandez et al. [213], Robust Multi-bit [214], Christ et al. [215], Kuditipudi et al. [216], Unigram Watermark [217], Private Watermark [218]   , 5tnode]] ]] ] ]

### v4

@@ 2 Generation
$\clubsuit$ Video. In the pursuit of high-quality video generation, recent research has also turned to diffusion models. Early work [187, 188] are based on DDPM [75] for video generation. Current research extends text-to-image (T2I) diffusion models to text-to-video (T2V) generation. Early T2V efforts include Meta’s Make-A-Video [189], which extends a diffusion-based T2I model using spatiotemporally factorized diffusion, and Emu Video [190], a simplified two-stage successor. Google’s Imagen Video [191], Stability AI’s Stable Video Diffusion [192], Runway’s GEN-2 [193], and Pika [194] also contributed video generation pipelines. More recently, the field has advanced rapidly with powerful models such as OpenAI’s Sora [195], Google DeepMind’s Veo2 [196]. as shown in Fig. 2 c).
LAIM Generated Datasets (Recommended). Though building upon the success of T2I generation, T2V generation requires temporally smooth and realistic motion that matches the text, besides high-quality visual content, making it still in the nascent stage compared with image generation. Furthermore, video generation is much more computationally costly than image synthesis. Recent work has begun to contribute to LAIM-generated video datasets, see Table I. DiffHead [118] provides high-quality talking-head videos generated from a single image and audio using a diffusion model. While domain-specific, its temporal smoothness and realistic expressions make it well-suited for evaluating detectors on subtle artifacts like lip-sync mismatch and unnatural facial motion. On the other hand, DVF [120] provides a broader testing ground for general-purpose video detection. It consists of fake videos generated by eight different diffusion models, including both open-source and commercial systems. GenVideo [123] and GenVidBench [127] represent the latest efforts toward scaling and realism. GenVideo provides an extensive corpus of over 2.3 million videos, collected from diverse real-world sources and spanning 20 generative models. Its emphasis on massive diversity and model heterogeneity makes it valuable for studying broad generalization across unseen generators. GenVidBench focuses on constructing a challenging benchmark for AI-generated video detection, emphasizing cross-source and cross-generator generalization.
Social Media. Deepfake videos are increasingly shared across platforms like YouTube, Twitter, and TikTok, highlighting the need for datasets that reflect the real-world complexity of such content. Unlike lab-created datasets—produced under controlled conditions with limited diversity—videos circulating online exhibit greater variability in quality, compression, and context. These differences are critical, as detectors trained on lab data often fail to generalize to the noisy, unpredictable nature of in-the-wild deepfakes seen on social platforms. WildDeepfake [117] collects over 700 deepfake videos directly from the internet, capturing a wide range of social media-style content with diverse backgrounds, multiple speakers, varying resolutions, and real-world compression effects, making it an effective benchmark for evaluating detector robustness in open-world scenarios. Complementing this, the Political Deepfakes Incidents Database (PDID) [130] compiles verified political deepfakes that have circulated widely on social platforms, enabling the study of how deepfakes influence discourse, political trust, and public perception. It supports detection research that goes beyond technical accuracy, addressing detection failures and social impacts.
Non-LAIM Generated Datasets. FaceForensics++ (FF++) [179], DeepfakeDetection (DFD) [180], Deepfake Detection Challenge (DFDC) [181], and Celeb-DF [182] are four most representative deepfake video datasets. Since they have also been used for image-level detection and were previously discussed, we omit their detailed discussion here to avoid redundancy.
@@
LAIM Generated Datasets (Recommended). Due to comparatively less research attention on the detection of fake audio, there are less LAIM-generated audio datasets compared to text and image. As shown in Table I, ASVspoof 2024 [140] introduces the most comprehensive benchmark to date, incorporating crowdsourced bona fide and spoofed speech from over 4,000 speakers and covering 32 manipulation methods, including advanced TTS ZMM-TTS [201], voice conversion (VC) DiffVC [202], and adversarial audio filtered through common codecs. This makes it particularly suitable for evaluating detection robustness in telephony and speaker verification scenarios. SONAR [136] focuses on evaluating generalization in audio deepfake detection, compiling over 2,200 AI-generated samples from nine state-of-the-art TTS models, including commercial systems like OpenAI [137] and Seed-TTS [138], which pose significant challenges for current detectors due to their high fidelity and proprietary training. Finally, LibriSeVoc [132] takes a focused approach by isolating neural vocoder artifacts through self-vocoding of real speech using six different vocoders, including DiffWave [134] and WaveGrad [198], enabling fine-grained artifact-based detection via a multi-task learning strategy.
Social Media. Müller et al. [131] introduce an in-the-wild audio deepfake dataset to reflect the challenges of detection in real-world, social media contexts. Unlike lab-controlled datasets like ASVspoof, their collection includes 37.9 hours of audio from 58 celebrities and politicians, sourced from social platforms and matched for speaker style, emotion, and noise. This dataset captures typical social media artifacts and content absurdity, making it highly realistic. Evaluation shows that models trained on ASVspoof 2019 [203] degrade by about 200% to 1000% EER when tested on this dataset, highlighting the urgent need for robust, socially grounded benchmarks for audio deepfake detection.
Non-LAIM Generated Datasets. Before the rise of LAIMs (e.g., diffusion models and LLMs), a number of influential audio datasets (e.g., WaveFake [204], ASVspoof 2019 [203], and ASVspoof 2021 [205]) were developed using traditional generative techniques such as GAN-based vocoders, like MelGAN [206], HiFiGAN [207]. Traditional spoofed speech is typically created from limited speaker data using purpose-built models with shallow linguistic understanding and observable signal-level artifacts. In contrast, LAIM-generated audio—such as from OpenAI [137], VALL-E [208], or Voicebox [209]—draws on massive multilingual corpora and produces highly coherent, context-aware, and personalized speech, often free of detectable synthesis artifacts. As such, detectors trained on non-LAIM data often fail to generalize to LAIM scenarios, underscoring the need for new benchmarks and detection methods specifically tailored to the emerging class of foundation model-based audio generation. WaveFake[204] comprises over 117,000 audio clips synthesized from six neural vocoders (e.g., MelGAN, HiFi-GAN), focusing on signal-level differences across architectures. Similarly, ASVspoof 2019[203] covers logical and physical access scenarios using VC and TTS techniques including Tacotron2 [210] and WaveNet [135]. The ASVspoof 2021 [205] marked a progression toward realism by introducing codec-transcoded, compressed, and replayed speech in semi-wild conditions, but still relied on predefined spoofing algorithms, not modern autoregressive or diffusion-based LLM architectures.
@@
$\clubsuit$ Multimodal. Multimodal learning refers to an embodied learning situation that includes learning multiple modalities such as text, image, video, and audio [211]. From a generation perspective, visual generation tasks, such as text-to-image and text-to-video, are regarded as multimodal generation. The generative models for these tasks are trained to learn visual representations and language understanding for visual generation. From the detection aspect, detectors that learn multiple modalities for forgery detection are categorized into multimodal. In this context, we define "multimodal generation" from a detection perspective, referring to frameworks capable of creating multimodal output. Multimodal generation process normally contains foundation models as encoders (e.g., CLIP [212], ViT [213] ) and decoders (e.g., Stable Diffusion [90] ), and a LLM for taking language-like representations from encoders for semantic understanding and produces modality signal tokens for guiding content output, see Fig. 2 e).
Most recent work, such as OpenAI GPT-4o [214], Google Gemini 2.5 [162], Meta LLaMA4 [215], Claude 3.5 Sonnet [216], DeepSeek V3 [217] , demonstrates rapid progress in unified multimodal modeling. These models are capable of processing and reasoning over diverse modalities with a single framework.
LAIM Generated Datasets (Recommended). The multimodal datasets listed in Table I capture diverse forms of synthetic media, spanning text, image, video, and audio. Early datasets such as $DGM^{4}$[14] and COCOFake[145] focused on image-text pair manipulations. $DGM^{4}$ introduces the task of detecting and grounding manipulations in both images and texts simultaneously, providing fine-grained annotations (e.g., bounding boxes, token-level edits) and simulating realistic multimodal misinformation scenarios such as fake news. COCOFake, in contrast, leverages COCO captions and Stable Diffusion to create over 1.2M text-guided fake images, highlighting both semantic and perceptual mismatches in multimodal generation. More recently, $M^{3}A$[152] and LOKI[154] have expanded multimodal detection benchmarks to unprecedented scale and modality diversity. $M^{3}A$ contains 6.5 million manipulated news samples across text, image, audio, and video, sourced from 60 reputable outlets and generated using diverse strategies like named entity manipulation and modality mismatching. It supports a wide range of tasks including deepfake detection, out-of-context analysis, and fact-checking. LOKI evaluates synthetic detection across 5 modalities and 26 subcategories, with rich, fine-grained anomaly annotations and explainability-focused benchmarks. It includes not only binary classification, but also reasoning-based tasks such as abnormal detail selection and explanation.
Social Media. MMFakeBench [155] advances the field with the first benchmark for mixed-source multimodal misinformation, combining three key forgery types: textual veracity distortion, visual veracity distortion, and cross-modal inconsistency. It contains 11K curated samples across 12 subcategories, revealing how current models struggle with realistic, multi-source misinformation. Deepfake-Eval [157] introduces an in-the-wild multimodal benchmark collected from social media and user submissions, covering video, audio, and image modalities, with over 2,400 generated samples and 3,300 real samples.
Non-LAIM Generated Datasets. In addition to LAIM-generated multimodal benchmarks, several earlier representative datasets were developed using non-foundational, task-specific generation methods. Notably, FakeAVCeleb[218] introduces a multimodal benchmark with paired deepfake videos and cloned audios, generated using models like FSGAN [219], Faceswap [220], and SV2TTS [221]. While it enables fine-grained evaluation across audio-video consistency, its generation methods rely on traditional GANs and vocoders, lacking the expressiveness and semantic generality of diffusion or autoregressive models. NewsCLIPpings[222] takes a different approach, focusing on out-of-context multimodal misinformation, where real images are automatically mismatched with real captions using CLIP-based semantic similarity. In contrast, recent LAIM-based datasets introduce greater content diversity, generative flexibility, and semantic complexity, making detection fundamentally more challenging and motivating the need for LAIM-focused detection research.
@@ 3 Detection
3.5.2)  , tnode, l sep=5mm, [$\clubsuit$ Empirical Study  , nnode,l sep=66.0mm, [Coccomini et al. [223],VERITE [224], Clipping [225], Jia et al. [226]   , 5mmnode]] [$\clubsuit$ Localization  , nnode,l sep=5mm, [ MLLM-based   , 4mmnode,l sep=5mm, [FakeShield [227], ForgeryGPT [228]   , 5mmnode]] [ Frequency-based   , 4mmnode,l sep=5mm, [UFAFormer [229]   , 5mmnode]] [ Spatial-based   , 4mmnode,l sep=5mm, [HAMMER [14], HAMMER++ [14], Wang et al. [230]   , 5mmnode]]] [$\clubsuit$ Interpretability  , nnode,l sep=66.0mm, [Xu et al. [231], FFAA [232], X²-DFD [233]   , 5mmnode]] [$\clubsuit$ Generalization  , nnode,l sep=5mm, [ Contrastive Learning   , 4mmnode,l sep=5mm, [LASTED [234]   , 5mmnode]] [ Prompt Tuning   , 4mmnode,l sep=5mm, [AntifakePrompt [235]   , 5mmnode]] ] [$\clubsuit$ Attribution  , nnode,l sep=66.0mm, [De-Fake [236], FIDAVL [237]   , 5mmnode]]] [Pure
3.5.1)  , tnode,l sep=5mm, [$\clubsuit$ Text-image Inconsistency  , nnode,l sep=66.0mm, [DIDAN [13], D-TIIL [238]   , 5mmnode]] [$\clubsuit$ Prompt-guided  , nnode,l sep=66.0mm, [Amoroso et al. [145], MM-Det [120], Human Action CLIPS [239]   , 5mmnode]] ] ] [Audio
3.4), xnode, l sep=5mm, [Beyond
3.4.2)  , tnode, l sep=5mm, [$\clubsuit$ Generalization  , nnode,l sep=66.0mm, [Ren et al. [240]   , 5anode]]] [Pure
3.4.1)  , tnode, l sep=5mm, [$\clubsuit$ Vocoder-based  , nnode,l sep=66.0mm, [DetectVocoder [132]   , 5anode]]] ] [Video
@@
3.3.2)  , tnode, l sep=5mm, [$\clubsuit$ Empirical Study  , nnode,l sep=66.0mm, [Vahdati et al. [241]   , 5vnode]] [$\clubsuit$ Generalization  , nnode,l sep=66.0mm, [RevisitVideo [242]   , 5vnode]]] [Pure
3.3.1)  , tnode, l sep=5mm, [$\clubsuit$ Spatial&Temporal-based Methods  , nnode,l sep=66.0mm, [DuB3D [243], He et al. [244]   , 5vnode]]] ] [Image
@@
3.2.2)  , tnode,l sep=5mm, [$\clubsuit$ Empirical Study  , nnode,l sep=66.0mm, [Corvi et al. [102, 245], Ricker et al. [97], Cocchi et al. [246], Papa et al. [71], Porcile et al. [247], Mandelli et al. [76], Carriere et al. [248]   , 5inode]] [$\clubsuit$ Robustness  , nnode,l sep=5mm, [ Post-Processing Robustness   , 4inode,l sep=5mm, [GLFF [249], Xu et al. [250], Local Statistics [251]   , 5inode]] [ Adversarial Attack Robustness   , 4inode,l sep=5mm, [D4 [252], De Rosa et al. [253]   , 5inode]]] [$\clubsuit$ Localization  , nnode,l sep=5mm, [ Weakly-supervised   , 4inode,l sep=5mm, [Tantaru et al. [254]   , 5inode]] [ Fully-supervised   , 4inode,l sep=5mm, [IFDL [92], PAL [255], TruFor [96], UnionFormer [256]   , 5inode]]] [$\clubsuit$ Interpretability  , nnode,l sep=66.0mm, [Aghasanli et al. [257], Color Statistics [258]   , 5inode]] [$\clubsuit$ Generalization  , nnode,l sep=66.0mm, [Online Detection [259], Ojha et al. [98], Cozzolino et al. [260], LSDA [261], Fingerprintnet [262], FACTOR [263], Dogoulis et al. [264], DNF [265] CoDE [151], HRR [266], AIDE [112], SUR-LID [267], B-Free [268], Breaking [269]   , 5inode]] [$\clubsuit$ Attribution and Model Parsing  , nnode,l sep=5mm, [ Model Parsing   , 4inode,l sep=5mm, [Asnani et al. [100]   , 5inode]] [ Attribution   , 4inode,l sep=5mm, [Guarnera et al. [270]   , 5inode]] ]] [Pure
3.2.1)  , tnode, l sep=5mm, [$\clubsuit$ Distribution-based Methods  , nnode,l sep=66.0mm, [ZED  [271]   , 5inode]] [$\clubsuit$ Frequency-based Methods  , nnode,l sep=66.0mm, [Wavelet-Packets [272], AUSOME [273], Xi et al. [274], Synthbuster [79], Lanzino et al. [275]   , 5inode]] [$\clubsuit$ Spatial-based Methods  , nnode,l sep=66.0mm, [AIGCD [103], Nguyen et al. [276], multiLID [277]   , 5inode]] [$\clubsuit$ Diffuser Fingerprints   , nnode,l sep=66.0mm, [DIF [104], DIRE [93], SeDID [278],LaRE² [279], Rajan et al. [280], Brokman et al. [281]   , 5inode]] [$\clubsuit$ Physical/Physiological  , nnode,l sep=66.0mm, [Borji [282], Perspective [283], Lighting [284]   , 5inode]] [$\clubsuit$ Watermarking  , nnode,l sep=66.0mm, [WaDiff [285], Stable Signature [286], Tree-Ring [287], Zhu et al. [288], Gaussian Shading [289]   , 5inode]] ] ] [Text
@@
3.1.2)  , tnode, l sep=5mm, [$\clubsuit$ Empirical Study  , nnode,l sep=66.0mm, [ChatLog [290], Zero-Shot [291], Training-based [48], Sarvazyan et al. [292], MAGE [34], Wahle et al. [30], Becker et al. [293], MGTBench [294], Guo et al. [50],Chen et al. [295],Antoun et al. [296]  , 5tnode]] [$\clubsuit$ Robustness  , nnode,l sep=5mm, [ LAIM-Polished Robustness  , 4tnode,l sep=5mm, [Yang et al. [46],   , 5tnode]] [ Adversarial Attack Robustness  , 4tnode,l sep=5mm, [Shi et al. [297], RADAR [298], OUTFOX [41], J-Guard [299]   , 5tnode]]] [$\clubsuit$ Interpretability  , nnode,l sep=66.0mm, [DNA-GPT [300],Kirchenbauer et al. [301], Mitrovic et al. [302], Check-GPT [60],Yang et al. [46]   , 5tnode]] [$\clubsuit$ Generalization  , nnode,l sep=66.0mm, [Ghostbuster [37],Conda [303],Text Fluoroscopy [304], DeTeCtive[305], PHD  [306]   , 5tnode]] [$\clubsuit$ Attribution  , nnode,l sep=66.0mm, [Linguistic Model [168], GPT-who [307],LLMDet [308], Few-Shot [309], Sniffer [310],TuringBench [311], Contra-X [312], XLNet-FT [313], TopFormer [314]   , 5tnode]] [$\clubsuit$ Efficiency  , nnode,l sep=66.0mm, [Efficient-DetectGPT [315], Fast-DetectGPT [316], DetectLLM [317], SeqXGPT [318], Glimpse [319]   , 5tnode]]] [Pure
3.1.1)  , tnode, l sep=5mm, [$\clubsuit$ Hard-explainable Methods  , nnode,l sep=66.0mm, [HowkGPT [320], GPTZero [321], Binoculars [322],DetectGPT [323],MPU [324]   , 5tnode]] [$\clubsuit$ Easy-explainable Methods  , nnode,l sep=5mm, [ Non-watermarking  , 4tnode,l sep=5mm, [Pu et al. [325],Stylometric Signals [326], Coherence Consistency [327], Zhu et al. [328], GECScore [329]   , 5tnode]] [ Watermarking  , 4tnode,l sep=5mm, [Distillation-Resistant [330], Fernandez et al. [331], Robust Multi-bit [332], Christ et al. [333], Kuditipudi et al. [334], Unigram Watermark [335], Private Watermark [336]   , 5tnode]] ]] ] ]